# Normalized AOPC: Fixing Misleading Faithfulness Metrics for Feature Attribution Explainability

### v1

@@ Discussion
Our findings demonstrate that AOPC scores must be normalized for cross-model comparisons and for meaningful interpretation of individual scores. For cross-model comparisons, normalization is necessary even when comparing models with identical architectures trained on the same dataset but with different seeds. Unnormalized scores can rank feature attribution methods for a single model, but they do not indicate how the scores relate to the model’s possible AOPC score range. Therefore, we recommend normalizing AOPC scores in all cases except when merely ranking attribution methods for a single model. These insights suggest that previous conclusions about model faithfulness may need re-evaluation, considering the importance of score normalization.
Why are some models more faithful than others?
We hypothesize that variations in model faithfulness stem from differences in how closely a model’s reasoning aligns with feature attribution methods’ assumptions. Faithfulness measures how accurately these methods’ assumed reasoning reflects the model’s actual process. Most feature attribution methods use simplified assumptions about model reasoning, such as feature independence (Bilodeau et al. 2024). Consequently, these methods may favor models processing features more independently (like linear models) over those using complex feature interactions. This bias is evident in our results, where models like RoBERTa consistently achieve low comprehensiveness scores across all tested attribution methods, even after normalization.
@@
In addition, we demonstrate that increasing NAOPC’s beam size from 5 to 50 had little effect on the results and, therefore, assumed that a beam size of 5 was sufficiently accurate. This assumption could be false. There could be a sudden jump in accuracy if the beam size were increased further. We stopped at beam size 50 because of our computational budget. However, its accuracy in examples with few features supports our claim of accuracy.
Related Work
Researchers have raised several criticisms against AOPC and other perturbation-based faithfulness metrics, which fall into three main categories. First, perturbing inputs can create out-of-distribution examples, potentially conflating distribution shifts with feature importance (Ancona et al. 2017; Hooker et al. 2019; Hase, Xie, and Bansal 2021). Second, perturbations often yield inputs that appear non-sensical to humans, though this should not affect faithfulness evaluation (Feng et al. 2018; Bastings and Filippova 2020; Jacovi and Goldberg 2020). Third, these metrics can be viewed as attribution methods themselves, potentially measuring similarity between methods rather than true faithfulness (Zhou and Shah 2022; Ju et al. 2023).
Alongside AOPC, the field has developed a wide range of faithfulness metrics. These generally fall into two categories: ordering-based and completeness-based. Ordering-based metrics measure whether attributions correctly order input features by importance. These include AOPC, Decision-flip metrics (Chrysostomou and Aletras 2022), Monotonicity and CORR (Arya et al. 2019), and soft-comprehensiveness and soft-sufficiency (Yin et al. 2021; Zhao and Aletras 2023). Completeness-based metrics assess how accurately each feature’s attribution score reflects its impact on the output, with sensitivity-n (Ancona et al. 2017) being the most adopted.
Conclusion
Our study exposes critical weaknesses in current faithfulness evaluation practices for feature attribution methods. We found that models’ reasoning processes significantly influence AOPC’s lower and upper limits, potentially leading to misleading cross-model comparisons. Moreover, without knowing these limits, it becomes difficult to interpret AOPC scores effectively. These findings challenge the validity of conclusions drawn from cross-model AOPC score comparisons in many influential studies on model faithfulness. To address these issues, we introduced NAOPC, a normalized measure that mitigates model-dependent bias while preserving the ability to compare feature attribution methods within individual models. NAOPC enables more accurate and reliable evaluations of feature attribution methods across different models, advancing the field towards more robust assessments of model faithfulness.
@@ Appendix A Model Details and Access
| Model | Architecture | Training data | HuggingFace link |
| BERT | BERT | Yelp | textattack/bert-base-uncased-yelp-polarity |
| RoBERTa | RoBERTa | Yelp | VictorSanh/roberta-base-finetuned-yelp-polarity |
| BERT | BERT | IMBD | textattack/bert-base-uncased-imdb |
| RoBERTa | RoBERTa | IMBD | textattack/roberta-base-imdb |
| BERT | BERT | SST2 | textattack/bert-base-uncased-SST-2 |
| RoBERTa | RoBERTa | SST2 | textattack/roberta-base-SST-2 |
@@ Appendix C NAOPC Comparison on SST2
Appendix D Impact of Beam Size on NAOPC Across Various Datasets
As illustrated in fig. 7, the relationship between increasing beam size and NAOPC is examined across various models and datasets. The findings indicate that while an initial expansion of beam size results in variability in the upper and lower bounds, further increases beyond a beam size of 5 lead to a convergence trend. This pattern is consistently observed across different models and datasets, particularly evident in the results for RoBERTa and BERT on the Yelp dataset.

### v2

@@ 6.2 Should one always normalize the AOPC scores?
For cross-model comparisons, normalization is necessary even when comparing models with identical architectures trained on the same dataset but with different random seeds, as even slight variations in model parameters can lead to different AOPC limits. Without normalization, a score of 0.25 could be near-optimal for one model but mediocre for another, making cross-model comparisons misleading. Therefore, we recommend normalizing AOPC scores in all cases except when only comparing the relative ranking of feature attribution methods within a single model, where the absolute values of the scores are not relevant. The fundamental importance of normalization for valid cross-model comparisons calls into question previous research findings. Studies that compared unnormalized AOPC scores across different models may need to be re-evaluated.
6.3 Why are some models more faithful than others?
By normalizing AOPC scores, we can now meaningfully compare explanation faithfulness across models and interpret how far they are from optimal performance. Our analysis reveals substantial differences in faithfulness even after normalization. For instance, on Yelp, all explanation methods achieved substantially higher comprehensiveness for BERT than RoBERTa. For DecompX, the best method, the difference was 0.75 and 0.44. Why do the explanation methods produce less faithful explanations for RoBERTa?
We hypothesize these variations stem from differences in how closely models align with the feature attribution methods’ assumptions. Most feature attribution methods rely on simplified assumptions about models’ inner mechanisms, such as feature independence Bilodeau et al. (2024). RoBERTa consistently achieves low comprehensiveness scores across all tested attribution methods, even after normalization, suggesting a fundamental mismatch between how this model processes information and current attribution methods’ assumptions. However, further research examining the models’ internal mechanisms would be needed to verify this hypothesis.
6.4 Why does normalization have a bigger impact on shorter text datasets?
Figure 2 and Figure 3 show that normalization changes the AOPC results more on shorter sequences than on longer. What causes this difference? We hypothesize that models typically rely on a small subset of features when processing long inputs. When measuring AOPC (comprehensiveness and sufficiency), we remove tokens based on their importance sequentially. In long texts, most removals have no effect since they were irrelevant in the model’s decision. When averaging across all removal steps, these zero-effect steps dilute the AOPC scores—bringing sufficiency scores closer to zero and comprehensiveness scores closer to one—making the impact of normalization less visible since both the raw scores and their theoretical limits converge. If there are fewer unused tokens in shorter texts, then fewer zero-effect steps are included in the average. This makes normalization effects more pronounced. If we could somehow consider only the tokens impacting the model’s decision, we hypothesize that the normalization effect would be similar on short and long inputs.
7 Related Work
Researchers have raised several criticisms against AOPC and other perturbation-based faithfulness metrics, which fall into three main categories. First, perturbing inputs can create out-of-distribution examples, potentially conflating distribution shifts with feature importance Ancona et al. (2017); Hooker et al. (2019); Hase et al. (2021). Second, perturbations often yield inputs that appear non-sensical to humans, though this should not affect faithfulness evaluation Feng et al. (2018); Bastings and Filippova (2020); Jacovi and Goldberg (2020). Third, these metrics can be viewed as attribution methods themselves, potentially measuring similarity between methods rather than true faithfulness Zhou and Shah (2022); Ju et al. (2023).
Our work focuses on sufficiency and comprehensiveness due to their widespread use in cross-model comparisons Bhalla et al. (2023); Li et al. (2023); Chrysostomou and Aletras (2022); Liu et al. (2022). However, researchers have also developed alternative faithfulness metrics to address the potential limitations of AOPC. Decision-flip metrics track when model predictions change as features are removed Chrysostomou and Aletras (2022). Monotonicity and the faitfulness correlation metric (CORR) measure whether higher attribution scores correspond to larger changes in model output Arya et al. (2019). Sensitivity-n tests if the sum of attribution scores equals the total change in model output when removing the features Ancona et al. (2017).
While these alternatives were designed to provide different perspectives on faithfulness, we suspect that they share some of AOPC’s fundamental limitations. Like AOPC, decision-flip metrics may produce misleading results when comparing models that rely on different numbers of features because fewer features need to be removed to significantly change the model’s prediction, resulting in artificially better scores. Similarly, we expect metrics like sensitivity-n, monotonicity, and CORR to struggle with feature interactions because they assume attribution scores can be assigned independently. This assumption likely breaks down when feature importance depends on feature interactions rather than independent features. For example, consider a model using OR operations (). Sensitivity-n requires attribution scores to sum to the total change when both features are removed () but also requires each score to equal its individual impact (), creating an impossible mathematical constraint. These limitations suggest a broader need to develop faithfulness metrics that can account for model-specific characteristics and complex feature interactions.
8 Conclusion
Our study exposes critical weaknesses in current faithfulness evaluation practices for feature attribution methods. Using simple toy models, we demonstrated how models’ inner mechanisms significantly influence AOPC’s lower and upper limits, potentially leading to misleading cross-model comparisons. Moreover, without knowing these limits, it becomes difficult to interpret AOPC scores effectively. These findings challenge the validity of conclusions drawn from cross-model AOPC score comparisons in many influential studies. To address these issues, we introduced NAOPC, a normalized measure that mitigates model-dependent bias while preserving the ability to compare feature attribution methods within individual models. NAOPC enables accurate evaluation of feature attribution faithfulness across different models, advancing the field towards more robust explanation assessment. While NAOPC addresses these fundamental issues, its computational complexity suggests the need for future research into faster interpretable faithfulness metrics that maintain cross-model comparability.
@@ 8 Conclusion
In addition, as discussed in Section 6.1, with its time complexity, NAOPC will be prohibitively slow for certain datasets, especially for those requiring large beam sizes. Most of our datasets and models required small beam sizes (B=5), but AG-News required a large beam size (B=1000). However, it is better with a slow evaluation than an inaccurate one. Moreover, we provide software tools to help researchers determine the necessary beam size for their specific use case. This allows researchers to assess the computational requirements beforehand and plan their experiments accordingly, deciding whether cross-model comparisons are feasible for their dataset and computational budget.
Ethical considerations
The ability to explain deep neural network decisions is crucial for ensuring their responsible deployment, particularly in high-stakes domains such as healthcare, legal systems, and financial services. When a diagnostic model suggests treatment or when a neural network influences a parole decision, stakeholders must be able to scrutinize and validate the reasoning behind these recommendations. However, explanations are only valuable if they faithfully represent the model’s decision-making process.
Our work reveals that current methods for evaluating explanation faithfulness can be misleading, potentially giving false confidence in explanation methods that do not accurately reflect model behavior. This is particularly concerning because unreliable explanations might lead to unwarranted trust in neural networks or mask potential biases in their decision-making processes. For instance, Kayser et al. (2024) demonstrated that incorrect explanations can persuade phycisians into an incorrect diagnosis. By providing a more reliable evaluation framework through NAOPC, we contribute to the development of more trustworthy explanation methods, ultimately supporting the responsible deployment of deep neural networks in society.
@@ Appendix A Model Details and Access
Model Architecture (Param) Training data HuggingFace link BERT BERT (110M) Yelp textattack/bert-base-uncased-yelp-polarity RoBERTa RoBERTa (110M) Yelp VictorSanh/roberta-base-finetuned-yelp-polarity BERT BERT (110M) IMBD textattack/bert-base-uncased-imdb RoBERTa RoBERTa (110M) IMBD textattack/roberta-base-imdb BERT BERT (110M) SST2 textattack/bert-base-uncased-SST-2 RoBERTa RoBERTa(110M) SST2 textattack/roberta-base-SST-2 BERT BERT (110M) AG-News textattack/bert-base-uncased-ag-news RoBERTa RoBERTa (110M) AG-News textattack/roberta-base-ag-news DistilBERT DistilBERT (66M) AG-News textattack/distilbert-base-uncased-ag-news BERT BERT (110M) SNLI textattack/bert-base-uncased-snli DistilBERT DistilBERT (66M) SNLI textattack/distilbert-base-cased-snli GPT-2 GPT-2 (124M) SNLI varun-v-rao/gpt2-snli-model1
@@ Appendix C NAOPC Comparison on SST2
Appendix D NAOPC results on AG-News and SNLI
Figure 7 depicts the difference between AOPC and NAOPC on AG-News and SNLI. We see similar results as in Figure 2. However, on SNLI, LIME and IG seem to be better than DecompX. Also, GPT-2’s sufficiency and comprehensiveness scores are similar. We speculate this is because we perturb by replacing with the end-of-sequence token (GPT-2 does not support mask tokens nor pad tokens). End-of-sequence tokens in the middle of a sentence may quickly make the input out-of-distribution, therefore changing the model’s output, even when perturbing unimportant features.
Appendix E Impact of Beam Size on NAOPC Across Various Datasets
As illustrated in Figure 8, the relationship between increasing beam size and NAOPC is examined across various models and datasets. The findings indicate that while an initial expansion of beam size results in variability in the upper and lower bounds, further increases beyond a beam size of 5 lead to a convergence trend. This pattern is consistently observed across different models and datasets, particularly evident in the results for RoBERTa and BERT on the Yelp dataset.
Appendix F Licences
SNLI uses a cc-by-sa-4.0 license. AG-News does not have a specific license, but the authors state that it should only be used for non-commercial purposes333http://groups.di.unipi.it/~gulli/AG_corpus_of_news_articles.html. SST-2 and IMDB are both created by StanfordNLP, who do not specify a license, but writes that you must cite their papers if using the dataset444https://ai.stanford.edu/~amaas/data/sentiment/555https://nlp.stanford.edu/sentiment/. We used Yelp from Huggingface666https://huggingface.co/datasets/fancyzhx/yelp_polarity. The original webpage with the Yelp dataset and License no longer exists. Considering that thousands of other papers use this dataset, it is most likely okay to use, but we cannot guarantee it since we could not find its license. The language models from Huggingface use an MIT license.